Mr.O'Leary is 4 times as old as his son. In 16 years he will be only twice as old. What are their ages. (Hint: If the son is x years old now, how old will be in another 16 years?)
Answer by checkley77(12844)

4x=x son=x thus mr=4*son=4x
In 16 years they will be:
4x+16=2(x+16)
4x+16=2x+32
4x-2x=32-16
2x=16
x=16/2
x=8 years old for the son.
Dad=4*8=32 years old.
Proof:
4*8+16=2(8+16)
32+16=2*24
48=48